40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13519     Entry time: Tue Jan 9 21:38:00 2018
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS recovery 
  • Aligned IFO to IR.
    • Ran dither alignment to maximize arm transmission.
    • Centered Oplev reflections onto their respective QPDs for ITMs, ETMs and BS, as DC alignment reference. Also updated all the DC alignment save/restore files with current alignment. 
  • Undid the first 5 bullets of elog13325. The AUX laser power monitor PD remains to be re-installed and re-integrated with the DAQ.
    • I stupidly did not refer to my previous elog of the changes made to the X end table, and so spent ages trying to convince Johannes that the X end green alignment had shifted, and turned out that the green locking wasn't going because of the 50ohm terminator added to the X end NPRO PZT input. I am sorry for the hours wasted sad
    • GTRY and GTRX at levels I am used to seeing (i.e. ~0.25 and ~0.5) now. I tweaked input pointing of green and also movable MM lenses at both ends to try and maximize this. 
    • Input green power into X arm after re-adjusting previously rotated HWP to ~100 degrees on the dial is ~2.2mW. Seems consistent with what I reported here.
    • Adjusted both GTR cameras on the PSL table to have the spots roughly centered on the monitors.
    • Will update shortly with measured OLTFs for both end PDH loops.
    • X end PDH seems to have UGF ~9kHz, Y end has ~4.5kHz. Phase margin ~60 degrees in both cases. Data + plotting code attached. During the measurement, GTRY ~0.22, GTRX~0.45.

Next, I will work on commissioning the BEAT MOUTH for ALS beat generation. 

Note: In the ~40mins that I've been typing out these elogs, the IR lock has been stable for both the X and Y arms. But the X green has dropped lock twice, and the Y green has been fluctuating rather more, but has mangaged to stay locked. I think the low frequency Y-arm GTRY fluctuations are correlated with the arm cavity alignment drifting around. But the frequent X arm green lock dropouts - not sure what's up with that. Need to look at IR arm control signals and ALS signals at lock drop times to see if there is some info there.

Attachment 1: GreenLockStability.png  41 kB  | Hide | Hide all
GreenLockStability.png
Attachment 2: ALS_OLTFs_20180109.pdf  157 kB  Uploaded Tue Jan 9 23:54:12 2018  | Hide | Hide all
ALS_OLTFs_20180109.pdf
Attachment 3: ALS_OLTF_data_20180109.tar.bz2  14 kB  Uploaded Tue Jan 9 23:57:50 2018
ELOG V3.1.3-